A car gravelling initially at 16m/s begins to accelerate at 4m/s . find the speed and distance covered at 3 minute


1
Expert's answer
2021-03-29T08:57:34-0400

The speed covered after 3 minutes (180 s) of acceleration at 4 m/s2 from 16 m/s:


"v=v_0+at=16+4·180=736\text{ m/s}."

The distance covered during the three minutes:


"d=v_0t+\frac{at^2}{2}=67680\text{ m}."

However, it is unclear whether we need to find the distance the car traveled during the third minute or during the three minutes. If during the third minute: find the distance covered during 2 minutes of such acceleration:


"d_2=v_0t_2+\frac{at_2^2}{2}=30720\text{ m}."

The distance covered during the third minute:


"d_3=d-d_2=36960\text{ m}."